# End-to-End Visual Autonomous Parking via Control-Aided Attention

### v1

@@ III-B1 Perception backbone
We adopt a ResNet-18 [42] encoder shared across six surround-view RGB images. The resulting features are fused into a unified bird’s-eye view (BEV) representation using the Lift-Splat-Shoot (LSS) framework [43]. To regularize the backbone and ensure the BEV features capture structural cues beyond control supervision, we employ auxiliary heads for semantic segmentation and depth estimation, ensuring the BEV features encode both geometric and semantic cues useful for downstream modules.
III-B2 Feature Fusion
In addition to the BEV feature extracted from multi-view images, ego-motion and parking goal information are required for downstream planning. Compared to previous end-to-end parking approaches [13], which embed the target slot implicitly into the BEV feature before tokenization, we explicitly tokenize all input information. The BEV features and ego-motion are encoded via CNN and MLP modules into compact representations, and the ego-target coordinates are explicitly tokenized. These representations are then concatenated to form the final fused feature:
$\mathbf{f}_{\text{BEV}},\mathbf{f}_{\text{ego}},\mathbf{f}_{\text{goal}}$ $=\mathrm{CNN}(\mathbf{I}_{1:N}),\mathrm{MLP}(\mathbf{e}),\mathrm{MLP}(\mathbf{g}),$ (1)
@@ III-B2 Feature Fusion
where $\mathbf{I}_{1:N}$ denotes the set of $N$ multi-view images, $\mathbf{e}$ represents the ego-motion vector, $\mathbf{g}$ denotes the target parking slot coordinates in the ego-vehicle frame, $\mathbf{f}_{\text{BEV}}$, $\mathbf{f}_{\text{ego}}$, and $\mathbf{f}_{\text{goal}}$ are their corresponding tokenized features, and $\mathbf{f}_{\text{fused}}$ is the concatenated fused representation.
Target tokenization module(TTM): We explicitly convert the target parking slot coordinates into a token via $\mathbf{f}_{\text{goal}}=\text{MLP}(\mathbf{g})$. This explicit encoding preserves geometric information about the goal more effectively, improving trajectory prediction and control performance. The effect of this design is further analyzed in Sec. IV-D.
@@ III-B3 Learnable Motion Prediction
We leverage the previous four frames, including control commands $c_{t-4:t-1}$ and vehicle states $s_{t-4:t-1}$, to estimate the current-step displacement relative to the last frame. Frame-wise features are encoded by a CNN and aggregated with an LSTM: $h_{t}=\mathrm{LSTM}\!\left(\{\mathrm{CNN}([c_{k},s_{k}])\}_{k=t-4}^{t-1}\right),$ where $h_{t}$ is the temporal latent state. Two MLP heads predict the mean and variance of the 2D displacement: $\bm{\mu}_{t}=W_{\mu}h_{t}+b_{\mu},\quad\log\bm{\sigma}^{2}_{t}=W_{\sigma}h_{t}+b_{\sigma}.$ The vehicle position is then updated as $(x_{t},y_{t})=(x_{t-1},y_{t-1})+\bm{\mu}_{t},$ which also refreshes the ego-frame target coordinates for planning.
During training, the predicted displacement distribution is supervised by the ground-truth displacement $\Delta\mathbf{x}_{t}^{\text{GT}}=(x_{t}^{\text{GT}}-x_{t-1}^{\text{GT}},\;y_{t}^{\text{GT}}-y_{t-1}^{\text{GT}})$. We minimize the negative log-likelihood (NLL) loss:
$\mathcal{L}_{\text{dyn}}=-\log\mathcal{N}\!\left(\Delta\mathbf{x}_{t}^{\text{GT}}\;\middle|\;\bm{\mu}_{t},\;\bm{\sigma}_{t}^{2}\right),$ (2)
where $\mathcal{N}(x\mid\mu,\sigma^{2})$ denotes the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
Remark. This module is used only at inference to propagate states, while training directly uses ground-truth poses without optimizing this component.
III-B4 Control-Aided Attention (CAA)
The Control-Aided Attention (CAA) is introduced such that the attention module is learned to focus on control-sensive areas such that the perception module can provide more reliable information for the planning and control modules. CAA module predicts, for each BEV location, the expected control gradient in a self-supervised manner. Given fused perception features $\mathbf{f}_{\text{fused}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$, it outputs a BEV-shaped attention map $\widehat{\nabla}_{f}\in\mathbb{R}^{X_{b}\times Y_{b}}$: $\widehat{\nabla}_{f}=\mathrm{MLP}(\mathbf{f}_{\text{fused}}),$ which refines the fused features via an element-wise product with channel-wise broadcasting:
$\mathbf{f}_{\text{CAA}}(c,i,j)=\mathbf{f}_{\text{fused}}(c,i,j)\cdot\widehat{\nabla}_{f}(i,j).$ (3)
III-B5 Control and Waypoint Prediction
Both control and waypoint prediction modules take the CAA-refined features $\mathbf{f}_{\text{CAA}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$ as input. The control prediction head follows the structure of E2EParking [13], producing discrete control tokens autoregressively.
The waypoint prediction head explicitly models short-horizon 2D vehicle positions. Using only the current $\mathbf{f}_{\text{CAA}}$, it outputs future waypoints for the next few steps: $\mathbf{w}_{t+1:t+4}=\mathrm{WaypointDecoder}(\mathbf{f}_{\text{CAA}}),$ where $\mathbf{w}_{t+1:t+4}$ are the predicted waypoints in ego-centric coordinates.
@@ III-C Loss Function
To achieve robust autonomous driving in complex urban scenes, CAA-Policy is trained under a unified multi-task framework that jointly optimizes planning, perception, and self-supervised attention alignment. Complementary objectives regularize each other: planning (control, waypoints) benefits from perception (segmentation, depth), while the Control-Aided Attention (CAA) loss aligns perception features with decision signals. The overall training objective integrates these components into a single loss:
$\mathcal{L}_{\text{total}}=\lambda_{\text{ctrl}}\mathcal{L}_{\text{ctrl}}+\lambda_{\text{wp}}\mathcal{L}_{\text{wp}}+\lambda_{\text{seg}}\mathcal{L}_{\text{seg}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{CAA}}\mathcal{L}_{\text{CAA}},$ (4)
where $\lambda$ is weighting coefficients balancing each task and each loss function is described as follows:
@@
At each time step $t$, the ground-truth continuous control command $\tilde{c}_{t}^{\text{GT}}=(\text{steering}_{t},\text{throttle}_{t},\text{brake}_{t})$ is uniformly quantized into a discrete token index $c_{t}^{\text{GT}}$. Let $\hat{c}_{t}$ be the predicted token distribution over the discrete control vocabulary. The control loss is then defined as:
$\mathcal{L}_{\text{ctrl}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{c}_{t+k},c_{t+k}^{\text{GT}}),$ (5)
where $c_{t+k}^{\text{GT}}$ denotes the ground-truth discrete control token at the $k$-th future step and $\hat{c}_{t+k}$ is the corresponding predicted probability distribution.
@@
At each time step $t$, the short-horizon future continuous waypoints $\tilde{\mathbf{w}}_{t+1:t+4}^{\text{GT}}=[(x_{t+1},y_{t+1}),\dots,(x_{t+4},y_{t+4})]$ are uniformly quantized into discrete token indices $\mathbf{w}_{t+1:t+4}^{\text{GT}}=[w_{t+1}^{\text{GT}},\dots,w_{t+4}^{\text{GT}}]$. Let $\hat{w}_{t+k}$ be the predicted token distribution for the $k$-th future waypoint. The waypoint loss is then defined as:
$\mathcal{L}_{\text{wp}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{w}_{t+k},w_{t+k}^{\text{GT}}),$ (6)
where $w_{t+k}^{\text{GT}}$ denotes the ground-truth discrete waypoint token after quantization, and $\hat{w}_{t+k}$ is the predicted probability distribution.
@@
Following E2EParking [13], we supervise BEV semantic segmentation with cross-entropy $\mathcal{L}_{\text{seg}}=\text{CE}(\hat{M},M)$ and depth prediction with binary cross-entropy $\mathcal{L}_{\text{depth}}=\tfrac{1}{N_{\text{f}}}\sum_{i\in\mathcal{FG}}\text{BCE}(\hat{D}_{i},D_{i}),$ where $\hat{M}$ and $\hat{D}$ are predicted BEV semantic and depth maps, $M$ and $D$ are ground-truth maps, and $N_{\mathcal{FG}}$ is the number of foreground pixels $\mathcal{FG}$ with valid depth.
@@
CAA module predicts a BEV-shaped attention map $A\in\mathbb{R}^{H\times W}$ that approximates the control sensitivity of the corresponding BEV location. Inspired by Grad-CAM [44], during training, the ground-truth control gradient is obtained via automatic differentiation: $\nabla_{f}^{\ast}=\partial\text{Control}/{\partial\mathbf{f}_{\text{fused}}}.$
@@
$\mathcal{L}_{\text{CAA}}=\left\|A-\nabla_{f}^{\ast}\right\|_{1}.$
@@ IV Experiments
Experiments are conducted in the CARLA 0.9.11 simulator (Town04-Opt), a parking lot scenario with 64 slots, following the extended environment of E2EParking [13]. Training and validation data are collected from successful parking trajectories generated by traditional modular pipelines with ground-truth BEV segmentation inputs, starting from random positions towards specified targets. For testing, we evaluate parking in unseen scenarios around novel parking slots. Model robustness is assessed using TSR, TFR, NTSR, NTFR, CR, and OR metrics, as reported in Tab. II.
TABLE II: Evaluation metrics for autonomous parking.
@@
TSR Probability that the ego vehicle successfully parks into the target slot. Success: horizontal distance <0.6<0.6 m, longitudinal distance <1<1 m, orientation error $10^{\circ}$.
TFR Probability that the ego vehicle parks into the target slot but with unacceptable error.
NTR Probability that the ego vehicle parks into a non-target slot.
CR Probability of parking collision.
TR Probability that the ego vehicle fails to park within a specified time.
@@ IV-A Implementation details
Most of our implementation follows E2EParking [13], including data collection, BEV feature size and the corresponding spatial resolution, backbone architecture, depth image parameters, encoder–decoder attention setup, and the optimization framework.
@@
For expert demonstrations, rather than using human driving as in [13], we employ trajectories generated by a modular Hybrid A* planner, with control signals obtained from a bicycle model. Only successful and smooth trajectories are manually selected for training.
@@ IV-B Baselines
We compare CAA-Policy against two baselines: E2E baseline [13] and Modular Hybrid A* baseline.
TABLE III: Comparison of parking performance across E2E baselines, ablations, the modular baseline, and the full model (Ours). All results are reported as $\mu\pm\sigma$ (mean $\pm$ standard deviation) in percentage over evaluation episodes. Metric definitions are provided in Table II.
@@
Hybrid A* 59.1 $\pm$ 3.8 5.5 $\pm$ 2.2 0.0 $\pm$ 0.0 12.2 $\pm$ 2.9 21.1 $\pm$ 3.7
E2EParking (Reproduced) [13] 26.8 $\pm$ 5.1 3.5 $\pm$ 1.7 19.9 $\pm$ 3.4 34.2 $\pm$ 5.2 10.4 $\pm$ 3.2
E2EParking (Original) [13] 74.7 $\pm$ 5.0 1.0 $\pm$ 0.6 3.1 $\pm$ 1.3 12.0 $\pm$ 3.8 8.9 $\pm$ 3.3
CAA-Policy (Ablation) 18.7 $\pm$ 4.4 2.0 $\pm$ 1.1 20.3 $\pm$ 3.4 32.8 $\pm$ 4.8 18.0 $\pm$ 4.0
@@ IV-B1 E2E baseline study
We first examine the E2EParking [13] baseline, an end-to-end model predicting control commands from raw multi-view images. While the original reported TSR is 74.7%, it was obtained with a substantially larger training set and extended training schedule , which are not available for our experiments. Using only 820 expert trajectories and half the training time, our reproduction achieves a TSR of 26.8% with a Collision Rate of 34.2%, highlighting the sensitivity of end-to-end approaches to dataset scale: under limited data, E2EParking fails to generalize to unseen parking slots and cannot reliably track the target.

### v2

@@ III-B1 Perception backbone
We adopt a ResNet-18 [he2016deep] encoder shared across six surround-view RGB images. The resulting features are fused into a unified bird’s-eye view (BEV) representation using the Lift-Splat-Shoot (LSS) framework [philion2020lift]. To regularize the backbone and ensure the BEV features capture structural cues beyond control supervision, we employ auxiliary heads for semantic segmentation and depth estimation, ensuring the BEV features encode both geometric and semantic cues useful for downstream modules.
III-B2 Feature Fusion with Target Tokenization Module (TTM)
In addition to the BEV feature extracted from multi-view images, vehicle state and parking goal information are required for downstream planning. Compared to previous end-to-end parking approaches [yang2024e2e], which embed the target slot implicitly into the BEV feature before tokenization, we explicitly tokenize all input information. The BEV features and vehicle states are encoded via CNN and MLP modules into compact representations, and the ego-target coordinates are explicitly tokenized. These representations are then concatenated to form the final fused feature:
$\mathbf{f}_{\text{BEV}},\mathbf{f}_{\text{ego}},\mathbf{f}_{\text{goal}}$ $=\mathrm{CNN}(\mathbf{I}_{1:N}),\mathrm{MLP}(\mathbf{s}),\mathrm{MLP}(\mathbf{g}),$ (1)
@@ III-B2 Feature Fusion with Target Tokenization Module (TTM)
where $\mathbf{I}_{1:N}$ denotes the set of $N$ multi-view images, $\mathbf{s}$ represents the vehicle state, $\mathbf{g}$ denotes the target parking slot coordinates in the ego-vehicle frame, $\mathbf{f}_{\text{BEV}}$, $\mathbf{f}_{\text{ego}}$, and $\mathbf{f}_{\text{goal}}$ are their corresponding tokenized features, and $\mathbf{f}_{\text{fused}}$ is the concatenated fused representation.
The encode the target effectively, we explicitly convert the target parking slot coordinates into a token via $\mathbf{f}_{\text{goal}}=\text{MLP}(\mathbf{g})$. This explicit encoding preserves geometric information about the goal more effectively, improving trajectory prediction and control performance. The effect of this design is further analyzed in Sec. IV-D.
@@ III-B3 Learnable Motion Prediction
We leverage the previous four steps of control commands $\mathbf{c}_{t-4:t-1}$ and vehicle states $\mathbf{s}_{t-4:t-1}$ to estimate the 2D displacement relative to the last frame. Frame-wise features are encoded by a CNN and aggregated by an LSTM into the temporal latent state $h_{t}=\mathrm{LSTM}(\{\mathrm{CNN}([\mathbf{c}_{k},\mathbf{s}_{k}])\}_{k=t-4}^{t-1})$. Two MLP heads predict the displacement mean and variance vectors $\bm{\mu}_{t}$ and $\bm{\sigma}_{t}^{2}$, and the vehicle position is updated as $(x_{t},y_{t})=(x_{t-1},y_{t-1})+\bm{\mu}_{t}$, which refreshes the ego-frame target coordinates. Here, $\bm{\mu}_{t}$ and $\bm{\sigma}_{t}^{2}$ represent per-axis displacement mean and variance.
During training, the predicted displacement $\bm{\mu}_{t}$ is supervised by the ground-truth displacement $\Delta\mathbf{p}_{t}^{\ast}=(x_{t}^{\ast}-x_{t-1}^{\ast},\;y_{t}^{\ast}-y_{t-1}^{\ast})$ using the NLL loss: $\mathcal{L}_{\text{dyn}}=-\sum_{i\in\{x,y\}}\log\mathcal{N}(\Delta p_{t,i}^{\ast}\mid\mu_{t,i},\sigma_{t,i}^{2})$, where each axis is treated independently.
Remark. This module is only used during inference, and itself is trained separately; during the end-to-end training of the major perception and planning system, ground-truth positions are directly used.
III-B4 Control and Waypoint Prediction
Both control and waypoint prediction modules take the CAA-refined features $\mathbf{f}_{\text{CAA}}\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}}$ as input. The control prediction head follows the structure of E2EParking [yang2024e2e], producing discrete control tokens autoregressively. Continuous control signals (steering, throttle, and brake) are discretized using the same scheme as in E2EParking, with identical bin boundaries, minimum, and maximum values for each control dimension.
The waypoint prediction head explicitly models short-horizon 2D vehicle positions. Using only the current $\mathbf{f}_{\text{CAA}}$, it outputs future waypoints for the next few steps: $\mathbf{w}_{t+1:t+4}=\mathrm{WaypointDecoder}(\mathbf{f}_{\text{CAA}}),$ where $\mathbf{w}_{t+1:t+4}$ are the predicted waypoints in ego-centric coordinates. The XY positions and heading angles are discretized using boundaries determined by statistical analysis of all training and testing trajectories: $\Delta x$ and $\Delta y$ are bounded within [-6, 6] meters, and the heading change $\Delta\psi$ is bounded within [-40, 40] degrees, with each dimension uniformly divided into 200 bins. These ego-centric waypoints are then transformed into global coordinates using the current vehicle pose at time $t$.
@@ III-B4 Control and Waypoint Prediction
III-B5 Control-Aided Attention (CAA)
The Control-Aided Attention (CAA) encourages the attention module to focus on control-sensitive areas, allowing the perception module to provide more reliable information for planning and control. For each BEV location, CAA predicts the expected control gradient in a self-supervised manner.
Given fused tokens $\mathbf{f}_{\text{fused}}$, we first encode them with a CNN to obtain a BEV feature map
$\mathbf{b}_{\text{BEV}}=\mathrm{CNN}(\mathbf{f}_{\text{fused}})\in\mathbb{R}^{C_{b}\times X_{b}\times Y_{b}},$
where $C_{b}$ is the number of channels and $X_{b},Y_{b}$ are the spatial dimensions of the BEV grid. CAA then outputs an attention map $A\in\mathbb{R}^{X_{b}\times Y_{b}},$ which refines the BEV features by spatially weighting each location (broadcasted across channels):
$\mathbf{f}_{\text{CAA}}[:,i,j]$ $=\mathbf{b}_{\text{BEV}}[:,i,j]\cdot A[i,j],$ (2)
$i=1,\dots,X_{b},\;j=1,\dots,Y_{b}.$
Gradients from the predicted control signals guide the attention, highlighting regions that most influence future actions and improving the robustness and generalization of downstream planning and control.
@@ III-C Loss Function
To achieve robust autonomous driving in complex urban scenes, CAA-Policy is trained under a unified multi-task framework that jointly optimizes control, waypoint prediction, perception, and self-supervised attention alignment. Complementary objectives regularize each other: control and waypoint prediction benefit from perception (segmentation and depth), while the Control-Aided Attention (CAA) loss aligns perception features with decision-making signals. The overall training objective integrates these components into a single loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{c}}+\mathcal{L}_{\text{w}}+\mathcal{L}_{\text{s}}+\mathcal{L}_{\text{d}}+\mathcal{L}_{\text{C}}.$ (3)
@@
at each time step $t$, the ground-truth continuous control command $\tilde{c}_{t}^{{}^{\ast}}=(\text{steering}_{t},\text{throttle}_{t},\text{brake}_{t})$ is uniformly quantized into a discrete token index $c_{t}^{\ast}$. Let $\hat{c}_{t}$ be the predicted token distribution over the discrete control vocabulary. The control loss is then defined as:
$\mathcal{L}_{\text{c}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{c}_{t+k},c_{t+k}^{\ast}),$ (4)
where $c_{t+k}^{\ast}$ denotes the ground-truth discrete control token at the $k$-th future step and $\hat{c}_{t+k}$ is the corresponding predicted probability distribution.
@@
at each time step $t$, the short-horizon future continuous waypoints $\tilde{\mathbf{w}}_{t+1:t+4}^{\ast}=[(x_{t+1},y_{t+1},\psi_{t+1}),\dots,(x_{t+4},y_{t+4},\psi_{t+4})]$ are uniformly quantized into discrete token indices $\mathbf{w}_{t+1:t+4}^{\ast}=[w_{t+1}^{\ast},\dots,w_{t+4}^{\ast}]$. Let $\hat{w}_{t+k}$ be the predicted token distribution for the $k$-th future waypoint. The waypoint loss is then defined as:
$\mathcal{L}_{\text{w}}=\sum_{k=1}^{4}\text{CrossEntropy}(\hat{w}_{t+k},w_{t+k}^{\ast}),$ (5)
where $w_{t+k}^{\ast}$ is the ground-truth discretized waypoint, and $\hat{w}_{t+k}$ the predicted probability distribution.
@@
Following E2EParking [yang2024e2e], we supervise BEV semantic segmentation with cross-entropy $\mathcal{L}_{\text{s}}=\text{CE}(\hat{M},M)$ and depth prediction with binary cross-entropy $\mathcal{L}_{\text{d}}=\tfrac{1}{N_{\text{f}}}\sum_{i\in\mathcal{FG}}\text{BCE}(\hat{D}_{i},D_{i}),$ where $\hat{M}$ and $\hat{D}$ are predicted BEV semantic and depth maps, $M$ and $D$ are ground-truth maps, and $N_{\mathcal{FG}}$ is the number of foreground pixels $\mathcal{FG}$ with valid depth.
@@
CAA module predicts a BEV-shaped attention map $A\in\mathbb{R}^{H\times W}$ that approximates the control sensitivity of the corresponding BEV location. Inspired by Grad-CAM [selvaraju2017gradcam], during training, the ground-truth control gradient is obtained via automatic differentiation: $\nabla_{f}^{\ast}=\partial\text{Control}/{\partial\mathbf{f}_{\text{fused}}}.$
@@
$\mathcal{L}_{\text{C}}=\left\|A-\nabla_{f}^{\ast}\right\|_{1}.$
@@ IV Experiments
Experiments are conducted in the CARLA 0.9.11 simulator (Town04-Opt), a parking lot scenario with 64 slots, following the extended environment of E2EParking [yang2024e2e]. Training and validation data are collected from successful parking trajectories generated by traditional modular pipelines with ground-truth BEV segmentation inputs, starting from random positions towards specified targets. For testing, we evaluate parking in unseen scenarios around novel parking slots. Model robustness is assessed using TSR, TFR, NTSR, CR and TR metrics, as reported in Tab. II.
TABLE II: Key evaluation metrics for autonomous parking.
@@
TSR Target Success Rate. Probability that the ego vehicle successfully parks into the target slot. Success: lateral distance <0.6<0.6 m, longitudinal distance <1<1 m, orientation error $10^{\circ}$.
TFR Target Failure Rate. Probability that the ego vehicle parks into the target slot but with unacceptable error.
NTSR Non-Target Success Rate. Probability that the ego vehicle parks into a non-target slot.
CR Collision Rate. Probability of collision with other obstacles.
TR Timeout Rate. Probability that the ego vehicle fails to park within 40 second.
@@ IV-A Implementation details
Most of our implementation follows E2EParking [yang2024e2e], including data collection, BEV feature size and the corresponding spatial resolution, backbone architecture, depth image parameters, encoder–decoder attention setup, and the optimization framework.
@@
For expert demonstrations, rather than using human driving as in [yang2024e2e], we employ trajectories generated by a modular Hybrid A* planner, with control signals obtained from a PID-controlled bicycle model. Only successful and smooth trajectories are manually selected for training.
@@ IV-B Baselines
We compare CAA-Policy against two baselines: E2E baseline [yang2024e2e] and Modular Hybrid A* baseline.
TABLE III: Comparison of parking performance across E2E baselines, ablations, the modular baseline, and the full model (Ours). All results are reported as $\mu\pm\sigma$ (mean $\pm$ standard deviation) in percentage over evaluation episodes. Metric definitions are provided in Table II; minor errors omitted.
@@
Hybrid A* [dolgov2008practical, kurzer2016path] 59.1 $\pm$ 3.8 5.5 $\pm$ 2.2 0.0 $\pm$ 0.0 12.2 $\pm$ 2.9 21.1 $\pm$ 3.7
E2EParking (Checkpoint) [yang2024e2e] 74.7 $\pm$ 5.0 1.0 $\pm$ 0.6 3.1 $\pm$ 1.3 12.0 $\pm$ 3.8 8.9 $\pm$ 3.3
E2EParking (Reproduced) [yang2024e2e] 17.6 $\pm$ 6.3 2.9 $\pm$ 1.7 29.4 $\pm$ 2.9 35.3 $\pm$ 5.5 14.7 $\pm$ 3.4
CAA-Policy (Ablation) 26.8 $\pm$ 5.1 3.5 $\pm$ 1.7 19.9 $\pm$ 3.4 34.2 $\pm$ 5.2 10.4 $\pm$ 3.2
@@ IV-B1 E2E baseline study
We begin with the E2EParking baseline [yang2024e2e], an end-to-end model that predicts control commands from raw multi-view images. Using the original checkpoint, it achieves a TSR of 74.7%, supported by a substantially larger training set and longer training schedule than those available in our setting. In contrast, our reproduction with only 820 expert trajectories and half the training time attains a TSR of 17.6% and a CR of 35.3%. These results highlight the sensitivity of end-to-end methods to dataset scale: with limited data, The baseline struggles to generalize to unseen parking slots and fails to reliably track the target.